OpenStudy (2yoututu11):

Which ratio is equivalent to 4 : 28? A. 1 : 8 B. 1 : 7 C. 8 : 52 D. 3 : 14

OpenStudy (anonymous):

b

OpenStudy (2yoututu11):



OpenStudy (ash2326):

\[\frac{4}{28}\] Divide numerator and denominator by 4 \[\huge \frac{\frac{4}{4}}{\frac{28}{4}}\] \[\huge \frac{\frac{\cancel 4 1}{\cancel 4}}{\frac{\cancel{28} 7}{\cancel 4}}\] We get \[\frac{1}{7}\]